OpenStudy (kaylala):

Find the value of the following using the concepts of both addition and subtraction formulas and reference angle: sin 255

OpenStudy (kaylala):

@MonoPolyRoly

OpenStudy (anonymous):

255 = 180 + ...?

OpenStudy (kaylala):

75

OpenStudy (anonymous):

So, you have sin(255) = sin(180+75) sin(180+x) = -sin(x) So, what would you get?

OpenStudy (kaylala):

how'd you do that????

OpenStudy (anonymous):

sin(180+75) = -sin(75)

OpenStudy (kaylala):

okay but i dont see any sin 75 in the graph

OpenStudy (anonymous):

75 = 45+...?

jhonyy9 (jhonyy9):

nice @MonoPolyRoly !!! good luck !

OpenStudy (kaylala):

45+75=75 @MonoPolyRoly what next?

OpenStudy (anonymous):

I am sure that 45+75 is not 75. If you find the correct answer for 75 = 45 + __, then you can apply the addition formula.